A solution of Initial Value Problem

(1-t2)(d2y/dt2) + 2t(dy/dt) - 2y = 0, y(0) = 3, y'(0) = -4

is y1=t. Use the method of reduction of order to find a general solution of Initial Value Problem on the interval -1<t<1.


1
Expert's answer
2021-06-03T15:26:45-0400

To solve

"(1-t^2)(\frac{d^2y}{dt^2}) + 2t(\frac{dy}{dt}) - 2y = 0"

such that

"y(0)=3, \text{ and } y'(0)=-4"

"\text{Let } y(t)=t v(t), \text{which gives }\frac{d y(t)}{d t}=t \frac{d v(t)}{d t}+v(t) \text{and }\frac{d^{2} y(t)}{d t^{2}}=t \frac{d^{2} v(t)}{d t^{2}}+2 \frac{d v(t)}{d t}\\
2 t\left(t \frac{d v(t)}{d t}+v(t)\right)+\left(-t^{2}+1\right)\left(t \frac{d^{2} v(t)}{d t^{2}}+2 \frac{d v(t)}{d t}\right)-2 t v(t)=0\\
\text{Simplify:}\\
-t \frac{d^{2} v(t)}{d t^{2}}\left(t^{2}-1\right)+2 \frac{d v(t)}{d t}=0\\
\text{Let }\frac{d V(t)}{d t}=u(t), \text{which gives }\frac{d^{2} v(t)}{d t^{2}}=\frac{d u(t)}{d t} :\\
\text{Solve for }\frac{d u(t)}{d t}:\\
\frac{d u(t)}{d t}=\frac{2 u(t)}{t\left(t^{2}-1\right)}\\
\text{Simplify:}\\
\frac{d u(t)}{d t}=-\frac{2 u(t)}{-t^{3}+t}\\"

"\text{Divide both sides by }u(t):\\
\frac{\frac{d u(t)}{d t}}{u(t)}=-\frac{2}{-t^{3}+t}\\
\text{Integrate both sides with respect to }t :\\
\int \frac{\frac{d u(t)}{d t}}{u(t)} d t=\int-\frac{2}{-t^{3}+t} d t\\
\text{Evaluate the integrals:}\\
\log (u(t))=-2\left(\log (t)-\frac{1}{2} \log \left(-t^{2}+1\right)\right)+c_{1}, \text{where }c_{1} \text{is an arbitrary constant.}\\
\text{Solve for }u(t) :\\
u(t)=-\frac{e^{c_{1}}\left(t^{2}-1\right)}{t^{2}}\\
\text{Simplify the arbitrary constants:}\\
u(t)=\frac{c_{1}\left(t^{2}-1\right)}{t^{2}}\\"

"\text{Substitute back for }\frac{d v(t)}{d t}=u(t) :\\
\frac{d v(t)}{d t}=\frac{c_{1}\left(t^{2}-1\right)}{t^{2}}\\
\text{Integrate both sides with respect to t :}\\
v(t)=\int \frac{c_{1}\left(t^{2}-1\right)}{t^{2}} d t=c_{1}\left(\frac{1}{t}+t\right)+c_{2}, \text{ where }c_{2} \text{ is an arbitrary constant.}\\
\text{Substitute back for }y(t)=t v(t), \text{ which gives }v(t)=\frac{y(t)}{t}:\frac{y(t)}{t}=c_{1}\left(\frac{1}{t}+t\right)+c_{2}\\
\text{Solve for }y(t):\\
y(t)=c_{1}+c_{1} t^{2}+c_{2} t\\
\text{Solve for the unknown constants using the initial conditions:}\\
\text{Compute }\frac{d y(t)}{d t}:\\
\frac{d y(t)}{d t}=\frac{d}{d t}\left(c_{1}+c_{1} t^{2}+c_{2} t\right)\\
=2 c_{1} t+c_{2}\\"

"\text{Substitute }y(0)=3 \text{ into }y(t)=c_{1}+t^{2} c_{1}+t c_{2}:\\
c_{1}=3\\
\text{Substitute }y^{\prime}(0)=-4 \text{ into }\frac{d y(t)}{d t}=2 t c_{1}+c_{2} :\\
c_{2}=-4\\
\text{Solve the system:}\\
c_{1}=3\\
c_{2}=-4\\
\text{Substitute }c_{1}=3 \text{ and }c_{2}=-4 \text{ into }y(t)=c_{1}+t^{2} c_{1}+t c_{2}:\\
\text{The required solution is:}\\
y(t)=3 t^{2}-4 t+3\\"